OpenStudy (anonymous):

Which of the following equations describe the line shown below? Check all that apply. (-3,-6) (1,6)

OpenStudy (eric_d):

where are the equations

OpenStudy (anonymous):

oh sorry eric

OpenStudy (anonymous):

@eric_d

OpenStudy (anonymous):

A. y-6=3(x-3) B. y+6=3(x+1) C. y-6=3(x+1) D. y+6=3(x+3) E. y=3x+2 F. y=3x+3

OpenStudy (anonymous):

@eric_d im done

OpenStudy (eric_d):

ok..

OpenStudy (anonymous):

check all that apply

OpenStudy (eric_d):

first ,find the gradient, m m= y2-y1/x2-x1= 3 So, rearrange the options into y=mX+C

OpenStudy (anonymous):

k den wat

OpenStudy (eric_d):

eliminate those options which don't hv gradient,m =3

OpenStudy (anonymous):

so a is out right

OpenStudy (eric_d):

yes

OpenStudy (anonymous):

k den b is out

OpenStudy (anonymous):

@eric_d is b out

OpenStudy (eric_d):

ys, u shld continue This method is easier y=mX+C Find C, y-intercept by subsituting value of x and y which is given.. like (1,6) y=mX+C 6=3(1)+C C=3 Final equation is y=3x+3 f.

OpenStudy (anonymous):

so the answers are c and f

OpenStudy (eric_d):

No.. it's f You think it's C coz I wrote there No! C means y-intercept

OpenStudy (anonymous):

okay so its just f im sorry im new to this im from germany

OpenStudy (eric_d):

No worries! You can use b instead of C as y-intercept

OpenStudy (anonymous):

i thought the answer was c i didnt think it would be f lol

OpenStudy (eric_d):

It is f I had shown u the working If u still don't understand, jst tag others... Bye.....

OpenStudy (anonymous):

k thanks